OpenStudy (anonymous):

Part 1: Solve the following system of equations. Please show your work to receive full credit. x - y = 19 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.

OpenStudy (anonymous):

Elimination: x-y=19 2x+y=2 + ------- 3x=21 x=7 Then substitute x into the first equation 7-y=19 y=-12 Check 2(7)+(-14)=12 It works

OpenStudy (anonymous):

thank you!